40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 67 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  13615   Wed Feb 7 15:50:42 2018 SteveUpdateVACIFO pressure monitoring

Hornet cold cathode gauge analoge output  [ DSub9 pin 3 and 7 ] are wired to go ETMX Acromag. It was reading 4.9V at 7.8e-6 Torr [ 3,110 V  8.35e-5A ] at the end of a 24ft BNC cable. Now it has to be hook up to an Acromag channel.

This will replace the not functioning C1: Vac-CC1_pressure

gautam: the motivation behind hooking this gauge up to our DAQ system is that non-vacuum-system-experts have a quick diagnostic to make sure everything is in order. This gauge is physically placed adjacent to V1, and so if something goes wrong with our vacuum pumps, we would see the effect here immediately. we did note that occassionally, the reading fluctuated by ~1V on the DMM used to check the voltage output at the end of the BNC cable, so we still need to run some long-term stability analysis once this channel is hooked up to the Acromag.  For future reference, in order to make this gauge work, we need to check that

  1. Error flag has been cleared.
  2. HV is ON (state has to be manually toggled).
Quote:

 

Quote:

There was a power outage.

The IFO pressure is 12.8 mTorr-it and it is not pumped. V1 is still closed. TP1 is not running. The Rga is not powered.

The PSL output shutter is still closed. 2W Innolight turned on and manual beam block placed in its beampath.

3 AC units turned on at room temp 84F

IFO pumped down from 44 mTorr to 9.6e-6 Torr with Maglev  backed with only TP3

Aux drypump  was helping our std drypump during this 1 hour period. TP3 reached 32 C and slowed down 47K rpm

The peak foreline pressure at P2  was ~3 Torr

Hornet cold cathode gauge setting:   research mode, air,

                                                            2830 HV  1e-4A  at 9.6e-6 Torr,

                                                         [  3110 HV  8e-5A at 7.4e-6 Torr one day later ]

Annuloses are at 2 Torr, not pumped

Valve configuration:  vacuum normal, RGA is still off

PSL shutter is opened automatically. Manual block removed.

End IR lasers and doublers are turned on.

 

NOTE: Maglev " rotation X " on vacuum medm screen is not working! " C1:Vac-TP1_rot " channel was removed.  Use " NORMAL X " for rotation monitoring.

*We removed this (i.e. rotation) field from the MEDM screen to avoid confusion.

 

Attachment 1: CC_analog.png
CC_analog.png
  13614   Wed Feb 7 12:35:56 2018 KiraUpdatePEMSeismometer can insulation test

I subtracted out the lab temperature change during the period of cooling to see if it would have a significant effect on the time constant, but when I fit the new data, the time constant came out to 0.355 hr, which is not a significant change from the value of 0.357 that I got earlier.

Attachment 1: fit_1.png
fit_1.png
  13613   Wed Feb 7 10:16:26 2018 gautamUpdateALSALS signal chain + power budget

After emailing the technical team at Menlo, I have uploaded the more detailed information they have given me on our wiki.

Quote:

The trouble is, I don't know what the transimpedance gain of the Fiber Beat PDs are. The datasheet suggests a "maximum gain" of 5e4 V/W, which presumably takes into account the InGaAs responsivity and the actual transimpedance gain.

 

  13612   Tue Feb 6 22:55:51 2018 gautamUpdateALSPossible source of ground loop identified

[koji, gautam]

We discussed possible solutions to this ground loop problem. Here's what we came up with:

  1. Option #1 - Configure the DAC card to receive a ground voltage reference from the same source as that which defines the LSC rack ground.
  2. Option #2 - construct an adapter that is differential-to-single ended receiving converter, which we can then tack on to these boards.
  3. Option #3 - use the D000186-revD board as the receiver for the DAC signals - this looks to have differential receiving of the DAC signals (see secret schematic).  We might want to modify the notches on these given the change in digital clock frequency 

Why do we care about this so much anyways? Koji pointed out that the tip tilt suspensions do have passive eddy current damping, but that presumably isn't very effective at frequencies in the 10Hz-1kHz range, which is where I observed the noise injection.

Note that all our SOS suspensions are also possibly being plagued by this problem - the AI board that receives signals is D000186, but not revision D I think. But perhaps for the SOS optics this isn't really a problem, as the expansion chassis and the coil driver electronics may share a common power source? 

gautam 1530 7 Feb: Judging by the footprint of the front panel connectors, I would say that the AI boards that receive signals from the DACs for our SOS suspended optics are of the Rev B variety, and so receive the DAC voltages single ended. Of course, the real test would be to look inside these boards. But they certainly look distinct from the black front panelled RevD variant linked above, which has differential inputs. Rev D uses OP27s, although rana mentioned that the LT1125 isn't the right choice and from what I remember, LT1125 is just Quad OP27...

  13611   Tue Feb 6 16:58:19 2018 KiraUpdatePEMtemperature measurements

I decided to plot the temperatures measured over two days for the sensor inside the can and inside the lab just to see if there was any significant difference between the two, and obtained the following plot. This shows that there is a difference in measurements of a few 0.01 C. The insulated seismometer can didn't change temperature as much as the lab did, which is as expected. I'll work on properly calibrating the sensors sometime in the future so that we can use the sensor that's just in the lab as an accurate thermometer.

Attachment 1: temps.png
temps.png
  13609   Tue Feb 6 11:13:26 2018 gautamUpdateALSPossible source of ground loop identified

I think I've narrowed down the source of this ground loop. It originates from the fact that the DAC from which the signals for this board are derived sits in an expansion chassis in 1Y3, whereas the LSC electronics are all in 1Y2.

  • I pulled the board out and looked at the output of Ch8 on the oscilloscope with the board powered by a bench power supply - signal looked clean, no evidence of the noisy ~20mVpp signal I mentioned in my previous elogyes.
  • Put the board back into a different slot in the eurocrate chassis and looked at the signal from Ch8  - looked clean, so the ground of the eurocrate box itself isn't to blameyes.
  • Put the board back in its original slot and looked at the signal from Ch8 - the same noisy signal of ~20mVpp I saw yesterday was evident again no.
  • Disconnected the backplane connector which routes signals from the DAC adaptor box to D000316 board - noisy signal vanished yes.

Looking at Jamie's old elog from the time when this infrastructure was installed, there is a remark that the signal didn't look too noisy - so either this is a new problem, or the characterization back then wasn't done in detail. The main reason why I think this is non-ideal is because the tip-tilt steering mirrors sending the beam into the IFO is controlled by analogous infrastructure - I confirmed using the LEMO monitor points on the D000316 that routes signals to TT1 and TT2 that they look similarly noisy (see e.g. Attachment #1). So we are injecting some amount (about 10% of the DC level) of beam jitter into the IFO because of this noisy signal - seems non-ideal. If I understand correctly, there is no damping loops on these suspensions which would suppress this injection. 

How should we go about eliminating this ground loop?

 

Quote:
 

So either something is busted on this board (power regulating capacitor perhaps?), or we have some kind of ground loop between electronics in the same chassis (despite the D990694 being differential input receiving). Seems like further investigation is needed. Note that the D000316 just two boards over in the same Eurocrate chassis is responsible for driving our input steering mirror Tip-Tilt suspensions. I wonder if that board too is suffering from a similarly noisy ground?

 

Attachment 1: A68AF89C-E8A9-416D-BBD2-A1AD0A51E0B5.jpeg
A68AF89C-E8A9-416D-BBD2-A1AD0A51E0B5.jpeg
  13608   Mon Feb 5 22:57:28 2018 gautamUpdateALSHuge harmonics in ALS channels

Did some quick additional checks to figure out what's going on here.

  1. The SOS/dewhite board for which I didn't have a DCC number is D000316. It has a single ended output.
  2. I confirmed that the origin of this noise has to do with the ground of the aforementioned D000316 - as mentioned in my previous elog, having one end of a BNC cable plugged into the whitening board D990694 and the other end terminated in 50ohms yields a clean spectrum. But making the ground of this terminator touch the ground of the SMA connector on the D000316 makes the harmonics show up.
  3. Confirmed that the problem exists when using either the SMA or the LEMO monitor output of these D000316 boards.

So either something is busted on this board (power regulating capacitor perhaps?), or we have some kind of ground loop between electronics in the same chassis (despite the D990694 being differential input receiving). Seems like further investigation is needed. Note that the D000316 just two boards over in the same Eurocrate chassis is responsible for driving our input steering mirror Tip-Tilt suspensions. I wonder if that board too is suffering from a similarly noisy ground?

Quote:
 

Am I missing something obvious here? I think it is impossible to do a THD measurement with the spectrum in this condition...

 

  13607   Mon Feb 5 18:04:35 2018 KojiUpdateComputer Scripts / Programsnetgpib data missing / PROLOGIX yellow box (crocetta) not working

crochetta was reconfigured to have 192.168.113.108. It was confirmed that it can be used with netgpibdata.py

Configuration: I have connected my mac with the unit using an Apple USB-Ethernet adapter. The adapter was configured to have a manual IP of 192.168.113.222/255.255.255.0. "netfinder.exe" was run to assign the IP addr to the unit. It seemed that NVRAM of the unit evaporated as it had the IP of 0.0.0.0. Once it was configrued, it could be run with netgpibdata as usual.

  13606   Mon Feb 5 14:11:01 2018 gautamUpdateALSHuge harmonics in ALS channels

I've been trying to setup for the THD measuremetn at the LSC rack for a couple of days now, but am plagued by a problem summarized in Attachment #1: there are huge harmonics present in the channel when I hook up the input to the whitening board D990694 to the output of a spare DAC channel at the LSC rack. Attachment #2 summarizes my setup. I've done the following checks in trying to debug this problem, but am no closer to solving it:

  • The problem is reminiscent of the one I experienced with the SR785 not too long ago - there the culprit was a switching power supply used for the Prologix GPIB-ethernet box.
  • Then I remembered sometime ago rana and i had identified the power supply for the Fibox at the LSC rack as a potential pollutant. But today, I confirmed that this power supply is not to blame, as I unplugged it from its powerstrip and the spectrum didn't change.
  • There are a couple of Sorensens in this LSC rack, from what it looks like, they supply power to a BIO interface box in the LSC rack. I thought we would want to keep this rack free of switching power supplies? Wasn't that the motivation behind keeping the (linear) power supplies for all the LSC rack electronics in a little separate rack along the east arm?
  • I confirmed that when the D990694 input is terminated, these harmonics are no longer present. 
  • I plugged the output of the SOS dewhite board to an oscilloscope - there is a ~20mVpp signal there even when the DAC output is set to 0, but this level seems too small to explain the ~1000 ct-pp signal that I was seeing. The whitening gains for these channels are set to 0dB.
  • I also looked at the signal in the time domain using DTT - indeed the peak-to-peak signal is a few thousand counts.
  • This isn't a problem with the particular input channel either - the behaviour can be reproduced with any of the 4 ALS input channels.

Am I missing something obvious here? I think it is impossible to do a THD measurement with the spectrum in this condition...

Attachment 1: ALSpowerlineHarmonics.pdf
ALSpowerlineHarmonics.pdf
Attachment 2: 2A1B3AC4-4059-416A-AD88-8AB0431D7A55.jpeg
2A1B3AC4-4059-416A-AD88-8AB0431D7A55.jpeg
  13605   Mon Feb 5 12:12:41 2018 KiraUpdatePEMSeismometer can insulation test

Attached the program I used to create the plot

Quote:

After taking the measurements, calibrating them (approximately), and filterting them, I created the following plot. The exponential fit is quite good, as the error is not more than 0.03 C. I used the python function curve_fit in order to get this, and it gave me the time constant as well, which came out to 0.357 hr. From my previous calculations here, I plugged in the values we have (m = 12.2 kg, c = 500 J/kg*k, d = 0.0762 m, k = 0.26 W/(m^2*K), A = 1 m^2), and got that

\tau = \frac{mcd}{kA}=0.496hr

This is a bit off, but it's probably due to the parameters not being exactly what I supposed them to be, and heat losses through the bottom of the can.

 

Attachment 1: temp_data.py
from scipy.optimize import curve_fit
from scipy.signal import decimate
import cdsutils as cds
import numpy as np
import matplotlib.pyplot as plt

#extract data
channel = ['C1:PEM-SEIS_EX_TEMP_OUT_DQ']
tstart = 1200962230
tend = 1201041084
... 49 more lines ...
  13604   Sat Feb 3 13:03:45 2018 gautamUpdateComputer Scripts / Programsnetgpib data missing / PROLOGIX yellow box (crocetta) not working

The netgpibdata scripts are now under git version control at /opt/rtcds/caltech/c1/scripts/general/labutils/netgpibdata. I think the idea was to make this directory a collection of useful utilities that we could then pull at various labs / at the sites.

Quote:

I could not understand why 'netgpibdata' scripts are missing in "scripts/general" folder on pianosa... Where did they go???

  13603   Fri Feb 2 23:28:13 2018 KojiUpdateComputer Scripts / Programsnetgpib data missing / PROLOGIX yellow box (crocetta) not working

I could not understand why 'netgpibdata' scripts are missing in "scripts/general" folder on pianosa... Where did they go???

Also, I found that the PROLOGIX GPIB-LAN controller for crocetta (192.168.113.108) is no longer working. I need to reconfigure it with "telnet"...

  13602   Fri Feb 2 22:47:00 2018 KojiSummaryGeneralAP1053: Packaging & Performance

I've packaged an AP1053 in a Thorlabs box. The gain and the input noise level were measured. It has the gain of ~10 and the input noise of ~0.6nV/rtHz@50MHz~200MHz.

Details

AP1053 was soldered on Thorlabs' PCB EEAPB1 (forgot to take a picture). The corresponding chassis is Thorlabs' EEA17. There is a 0.1uF high-K ceramic cap between DC and GND pins. The power is supplied via a DC feedthru capacitor (Newark / Power Line Filter / 90F2268 / 5500pF) found in the WB EE shop. The power cable has a connector to make the long side of the wires detachable. Because I did not want to leave the RF signal path just mechanically touched, the SMA connectors were soldered to the PCB. As the housing has no access hole, I had to make it at one of the sides.

The gain of the unit was measured using the setup shown in the upper figure of Attachment 2. When the unit was energized, it drew the current of about 0.1A. The measued gain was compensated by the pick off ratio of the coupler (20dB). The gain was measured with the input power of -20, -10, 0, 10, and 15dBm. The measurement  result is shown in Attachment 3. The small signal gain was actually 10dB and showed slight degradation above 100MHz. At the input of 10dB some compression of the gain is already visible. It looks consistent with the specification of +26.0dBm output for 1dB compression above 50MHz and +24.0dBm output below 50MHz.

The noise level was characterized with the setup shown in the bottom figure of Attachment 3. The noise figure of the amplifier is supposed to be 1.5dB above 200MHz and 3.5dB below 200MHz. This is quite low and the output noise of AP1053 can not be measured directly by the analyzer. So, another LN amplifier (ZFL-500HLN) was stacked. The total gain of the system was measured in the same way as above. The measured noise level was ~0.7nV/rtHz between 50MHz and 200MHz. Considering the measurement noise level of the system, it is consistent with the input referred noise of 0.6nV/rtHz. I could not confirm the advertized noise figure of 1.5dB above 200MHz. The noise goes up below 50MHz. But still 2nV/rtHz at 3MHz. I'd say this is a very good performance.

Attachment 1: AP1053.JPG
AP1053.JPG
Attachment 2: AP1053_measurement.pdf
AP1053_measurement.pdf
Attachment 3: AP1053_gain.pdf
AP1053_gain.pdf
Attachment 4: AP1053_noise.pdf
AP1053_noise.pdf
  13601   Fri Feb 2 21:12:46 2018 Udit KhandelwalSummaryGeneralSummary - 2018/02/02

Tip-Tilt Suspension CAD:
Discussed with Koji about motivation to simplify the design of this assembly, which has many unnecessary over-constraints. I have started to cad alternate parts with the aim of removing these over-constraints.

40m Lab CAD:
Acquired a stack of original engineering drawings of the vacuum chambers from Steve which I will take home, get scanned, and then use as reference for the cad i'm working on.

Other:
Started paperwork at west bridge office to get paid as an "occasional employee". Hopefully I receive old money.

  13600   Fri Feb 2 13:16:55 2018 gautamUpdateALSALS signals whitening switching

While setting up for this measurement, I noticed something odd with the whitening switching for the ALS channels. For the usual LSC channels, the whitening is set up such that switching FM1 on the MEDM screen changes a BIO bit which then enables/disables the analog whitening stage. But this feature doesn't seem to be working for the ALS channels - I terminated all 4 channels at the LSC rack, and measured the spectrum of the IN1 signals with DTT in the two settings, such that I expect to see a difference in the spectra if the whitening is enabled or disabled - FM1 enabled (expected analog whitening to be engaged) and FM1 disabled (expected analog whitening to be bypassed). But I see no difference in the spectra. I confirmed that the BIO bit switching is happening at least on the software level (i.e. the bit indicator MEDM screens indicate state toggling when FM1 is ON/OFF). But I don't know if something is amiss in the signal chain, especially since we are using Hardware channels that were previously used for AS_165 and POP_55 signals.

Is the whitening shape such that we expect the terminated noise level to be below ADC Noise even when the whitening is engaged? I just checked the shape of the de-whitening filter, and it has -40dB gain above 150Hz, so the inverse shape should have +40dB gain. 

Quote:
 

I will now proceed to the next piece (#3?) of this puzzle, which is to understand how the D990694 which receives the signals from this unit reacts to the expected DC voltage level of ~4Vpp


gautam 2.15pm: This was a FALSE ALARM, with the inputs terminated, the electronics noise really is that low such that it is buried under ADC noise even with +40dB gain. I cranked up the flat whitening gain from 0dB to 45dB for the X channels (but left the Y channels at 0dB). Attachment #2 is the comparison. Looks like the switching works just fine.

Attachment 1: ALS_whitening_switching.pdf
ALS_whitening_switching.pdf
Attachment 2: ALS_whitening_switching_works.pdf
ALS_whitening_switching_works.pdf
  13599   Fri Feb 2 00:26:34 2018 gautamUpdateALSD0902745 revamp underway

I saw some interesting behaviour of the Audio IF amplifier stage on the demod board today, by accident. I was testing the board for I/Q orthogonality and gain balance, when I noticed a large gain imbalance between the I and Q channels for both Board #3 and #4, which are the ones we use for the IR ALS demodulation. This puzzled me for some time, but then I realized that I had only reduced the gain of this stage from x100 to x10 for the I channel, and not for the Q channel! The surprising thing though was that the output waveform still looked like a clean sinusoid on the o'scope, and there was no evidence of the voltage clipping that is characteristic of an op-amp being driven beyond its voltage rails. The conversion factor with a preamp gain on x10 was measured today to be 2V IF / 1V RF. But this means that for a preamp stage gain of x100, we expect 20V IF / 1V RF, which is well in the saturation regime of the AD829, since the Vcc is only +/-15V. I'm guessing the diodes D2 and D3 are for overvoltage protection, but given that the pre-amp gain is x100, the input signal at the inverting input of the AD829 is only 0.2V at DC, which isn't above the forward bias voltage for the switching diode BAV99. Perhaps there is some interaction between the pre-amp and the FET demodulator that I dont understand, or I am missing something about the differential to single-ended topology that would explain this behaviour. 

I found it puzzling why the large preamp stage gain didn't hurt us with the green beat - even though the green optical beat signal was smaller than the current IR beat, a back-of-the-envelope calculation suggested that it would still have saturated the ADC with a x100 gain on the preamp. Perhaps this observation is part of the story, and there is also the unpredictable behaviour of the D990694 board for an input signal with large DC levels...


I did the following tests on this board today:

  • Check +/-15V supplies, power reg board.
  • Check DC offset on I and Q front panel output with LO driven at +10dBm, RF input terminated. Found it to be 0.
  • Checked calibration of back-panel DSUB connector monitors for LO and RF powers. Data to be uploaded, looked quite linear.
  • Checked conversion gain from RF input to IF output for two sets of LO/RF powers.
  • Measured conversion gain as a function of the IF frequency (i.e. frequency offset between LO and RF inputs, out to ~700kHz, 8 datapoints)
  • Checked orthogonality and gain balance of the I and Q outputs.
  • Measured the noise of the I and Q outputs in the audio frequency range using the SR785.

I didn't really measure the transfer function of the preamp stage after the modification because there wasn't a convenient test point and I couldn't find the high impedance FET probe for the Agilent - I wonder if somebody in WB has it? Anyways, all the tests suggested the board is operating as expected, and I now have calibrations for the back panel DSUB for LO/RF power levels, and also the conversion gain from RF to IF. I will put together a python notebook with all my measurements and upload it to the DCC page for this part. I need to double check expected noise levels from LISO to match up to the measurement.

I will now proceed to the next piece (#3?) of this puzzle, which is to understand how the D990694 which receives the signals from this unit reacts to the expected DC voltage level of ~4Vpp.


After discussion with Koji, I have also decided to look into putting together a daughter board for an alternative Audio IF preamp stage. The motivation is that for the ALS application, we expect a high DC signal level all the time (because the loop does not suppress the beat note amplitude). So we would like for the preamp stage to have the usual shape of some zero around 4Hz, a pole around 40Hz, and then the LowPass profile of the existing preamp stage (to cut out the 2f frequency product, but also to minimize the possibility of the fast AD829 going into some unpredictable regime where it oscillates). So, the desired features are:

  • Whitening (z,p) at (4Hz,40Hz) or (15Hz,150Hz) so that we have frequency dependent gain that can handle the large DC signal level expected. Need to measure noise of the actual IR beat signal to determine what the appropriate whitening shape is.
  • Low-pass above a few 100kHz to cut out 2f modulation product
  • Low-passing at input of AD829 (or just use OP27?)
  13598   Thu Feb 1 16:09:13 2018 KiraUpdatePEMSeismometer can insulation test

After taking the measurements, calibrating them (approximately), and filterting them, I created the following plot. The exponential fit is quite good, as the error is not more than 0.03 C. I used the python function curve_fit in order to get this, and it gave me the time constant as well, which came out to 0.357 hr. From my previous calculations here, I plugged in the values we have (m = 12.2 kg, c = 500 J/kg*k, d = 0.0762 m, k = 0.26 W/(m^2*K), A = 1 m^2), and got that

\tau = \frac{mcd}{kA}=0.496hr

This is a bit off, but it's probably due to the parameters not being exactly what I supposed them to be, and heat losses through the bottom of the can.

Attachment 1: cooling_fit.png
cooling_fit.png
  13597   Thu Feb 1 15:31:12 2018 gautamUpdateALSALS signal chain + power budget

Summary:

A reasonable level of RF beatnote power for operating within the specs of the demod board is 17dBm arriving at the input to the power splitter just before the delay line.

Details:

Stuff is beginning to look clearer now that I've done some initial characterization of the demod boards. I will upload a more detailed report of the characterization on the DCC page, but important findings are:

  1. The overall conversion factor from RF to IF is ~2.3V IF per volt of RF.
    • 50ohm source connected to RF input of demod board, level = 10dBm on Marconi screen, consistent with inferred value from RF mon output.
    • LO driven at 14dBm by Fluke function generator.
    • The ratio was calculated for IF voltage input into a High-Z load.
    • So let's say we want to run at half the ADC full range of 10Vpp into the whitening board - this means we need to keep the RF input to <=11dBm. 
  2. The Teledyne amplifier has a rated maximum input voltage of 17dBm. If we want to stay 3dB below this, we can send in 14dBm into the LO input of the demod board, which is what my characterizations were done with.

The delay line has a loss of ~3dB. The power splitter has a loss of 3dB. So putting everything together, 17dBm at the input of the power splitter gives us just the right amount of RF power to have the LO input driven at 14dBm, and the IF output be ~5Vpp into a High-Z load, which is about half the ADC full range.

 

  13596   Thu Feb 1 01:24:56 2018 gautamUpdateALSD0902745 revamp underway

I effected the change to the Audio IF preamp stage on channels 3 and 4 (Xarm and Yarm respectively) using the resistors Steve ordered (the ones Rana ordered don't have any labeling on them, and I couldn't tell the 50ohm and 500ohm ones apart except by looking at the label on the ziplock bag they came infrown, so I decided against using them). I've started a DCC page to collect photos, characterization data, and marked up schematic etc for this part. Characterization is ongoing, more to follow soon. Note that for the photo-taking, I disconnected all the on-board SMA connectors so that the cabling wouldn't block components. I have since restored them for testing purposes, and was careful to use the torque-limited SMA tightening tool when restoring the connections.

In order to test various things like conversion loss etc, I figured it would be useful to have two RF signal sources, so I scavenged the Fluke RF generator that Johannes was using from under the PSL table. In the process, I accidentally bumped the PSL interlock on the southeast corner of the PSL table. I immediately turned the NPRO back on, and relocked PMC/IMC. Everything looks normal now. Acromag may even have caught my transgression.

Quote:
 

I am going to characterize the demod board using E1100114. I am unsure as to the conversion loss of the mixer - the datasheet suggested a number of 8dB, but T1000044 suggests that the conversion loss is actually only 4dB. I figure it's best to just measure it. Would also be good to verify that the overall transfer function and noise of the IF amplifier stage match my expectation from the LISO model.

 

Attachment 1: PSLinterlock.png
PSLinterlock.png
  13595   Wed Jan 31 22:32:11 2018 gautamUpdateALSALS signal chain + power budget

Summary:

I do not have an answer to the question "What is an appropriate gain for the IF amplifier stage in the D0902745 FET demod boards?", because of the following problems.

Deatils:

The plan is to lower the gain of the IF amplifier stage on the FET demodulator board from 100 to 10. As per Attachment #1, this will make the overall gain from RF beatnote from the Beat Mouth to the signal input to the D990694 whitening board +19dB, assuming "typical" values for the conversion loss of the mixer, and the various other passive components on the FET demod board. I've used numbers I measured a couple of weeks ago for the delay line loss and the cabling loss from the PSL table to the LSC rack. This in turn will set a limit on how much RF beat power we can handle, from the Beat Mouth. According to this power budget, if we have -5dBm of beat, we will have an input to the whitening board of ~6Vpp, which is about half its full range. The trouble is, I don't know what the transimpedance gain of the Fiber Beat PDs are. The datasheet suggests a "maximum gain" of 5e4 V/W, which presumably takes into account the InGaAs responsivity and the actual transimpedance gain. However, according to the last power budget I did inside the Beat Mouth, I had -8dBm of beat for a combined 400uW of PSL+EX light, which definitely does not add up. I've emailed the company to ask about the spec, haven't gotten anything useful yet...

The problem is further complicated by the fact that the fiber inside the Beat Mouth is NOT polarization maintaining, and so the actual relative polarizations of the arm IR light and the PSL IR light is unpredictable, and also uncontrolled. I suppose we could simply place a HWP before the fiber collimator at either end, and rotate the polarization until we get a desired amount of beat, but this still does not solve the problem of the polarization being uncontrolled.

I am going to characterize the demod board using E1100114. I am unsure as to the conversion loss of the mixer - the datasheet suggested a number of 8dB, but T1000044 suggests that the conversion loss is actually only 4dB. I figure it's best to just measure it. Would also be good to verify that the overall transfer function and noise of the IF amplifier stage match my expectation from the LISO model.

Option #1: Rana ordered 50ohm and 500ohm SMD resistors of the 0805 package size, I asked Steve to get a few more values just in case we want to twiddle with the gain of this stage further (specifically, I asked for values such that we can set it to x5, x3 and x1). But changing the feedback resistors modifies the overall TF shape - see e.g. Attachment #2. Need to also look at how the noise performance varies.

Another possibility is to turn down the gain of the IF amplifier stage to x10, retire the ZHL-3A, and use a lower gain amplifier in its place. We do have the recently acquired Teledyne amplifiers, but we would have to package it in such a way that it can be integrated into the existing Fiber ALS signal chain. This would allow us to handle significantly larger RF beatnote powers, which I expect we will have if we improve the mode matching into the fibers (provided the aforementioned polarization drift possibility doesn't hurt us too much).

A third possibility is to attenuate the power coupled into the fibers to lower the RF beatnote amplitude. I don't like this option so much because placing an ND filter or a PBS+HWP combo in the beam path is likely to screw up the mode-matching into the fiber collimator, which I have already spent so many hours trying to improve, but if it must be done, it must be done.

The correct option is of course the one that gives us the lowest ALS noise. It is not clear to me which one that is at this point.

Attachment 1: FiberALS_PowerBudget.pdf
FiberALS_PowerBudget.pdf
Attachment 2: preampProposed.pdf
preampProposed.pdf
  13594   Wed Jan 31 16:33:53 2018 gautamUpdateALSALS electronics at LSC rack

[steve, gautam]

We installed some rails to mount the 2U chassis containing ~100m of delay line cabling, and the 1U chassis containing the FET demodulators for the ALS signals in the LSC rack. This has made it MUCH easier for a single person to work there and remove/reinstall these chassis. The delay line box has 100m of cable inside it, and so was rather heavy (~8kg) - previously, it was being supported only by a pair of brackets on the front, so the new arrangement is much more robustyes. Steve is looking into acquiring plastic spacers of the appropriate width, so that we can secure the units to the rack using usual rack mount screws (but the material of the newly installed rails and the screw heads holding them in place necessitate this plastic spacer). 

Delay line box has been re-installed, demodulator chassis has been removed by me for characterization. Steve will put up photos once the units are re-installed.

For this work, I had to disconnect a bunch of cabling, but only those connected to ALS. All cables were labelled, and I will re-connect them once I am done with the demod chassis.

Quote:
 

Anyways I think this is a big improvement from what was the situation before, and I am moving on to debugging the ALS electronics.

 

  13593   Wed Jan 31 16:29:42 2018 gautamUpdateALSModulation depths

I used the Beat Mouth to make a quick measurement of the PMC and EX modulation depths. They are, respectively, 60mrad and 90mrad. See Attachments #1 and #2 for spectra from the beat photodiode outputs, monitored using the Agilent analyzer, 16 averages, IF bandwidth set to resolve peaks offset from the main beat frequency peak by 33.5MHz for the PMC and by ~230kHz for the EX green PDH.

For this work, I had to re-align the IFO so as to lock the arms to IR. c1susaux was unresponsive and had to be power-cycled. As mentioned in the earlier elog, to avoid saturating the Fiber Coupled beat PDs, I placed a ND=0.5 filter in the fiber collimator path, such that the coupled power was ~1mW, which is well inside the safe regime.

For the EX modulation depth, I could have gotten multiple estimates of the modulation depth using the higher order products that are visible in the spectrum, but I didn't.

Attachment 1: PMCmodDepth.pdf
PMCmodDepth.pdf
Attachment 2: XPDH.pdf
XPDH.pdf
  13592   Wed Jan 31 15:46:05 2018 SteveUpdatesafetycrane inspection

Annual crane inspection with load tests is scheduled for Monday, Feb 5, 2018 from 8 to 11:30am

Konecranes rescheduled this appointment to: Monday, Feb 12, 2018

  13591   Wed Jan 31 15:45:22 2018 gautamUpdateALSFiber ALS assay

Attachment #1 shows the current situation of the PSL table IR pickoff. It isn't the greatest photo but it's hard to get a good one of this setup. Now there is no need to open the Green PSL shutter for there to be an IR beat note.

  • The key to improving the mode-matching was to abandon my "measurements" of the input mode and the mode from the collimator.
  • The best I could do with these measurements was ~25% coupling, whereas now I have ~78%yes (all powers measured with Ophir power meter).
  • Focusing was done using two f=300mm lenses (see attachment).
  • By moving the second (closer to collimator) lens through ~1inch of its current position, I was able to see a clear maximum of the coupled power.
  • By moving the second lens by ~5mm, and touching up the alignment, I couldn't see any improvement.

All this lead me to conclude that I have reached at least some sort of local maximum. The AR coating of the lens has ~0.5% reflection at 8 degrees AOI according to spec, and EricG mentioned today that the fiber itself probably has ~4% reflection at the interface due to there not being any special AR coating. There is also the fact that the mode of the collimator isn't exactly Gaussian. Anyways I think this is a big improvement from what was the situation before, and I am moving on to debugging the ALS electronics.

There is 3.65mW of power coupled into the fiber - our fiber coupled PDs have a damage threshold of 2mW, and this 3.65mW does get split by 4 before reaching the PDs, but good to keep this number in mind. For a quick measurement of the PMC and X end PDH modulation depth measurements, I used an ND=0.5 filter in the beam path.

 

Attachment 1: IMG_6875.JPG
IMG_6875.JPG
  13590   Wed Jan 31 15:29:44 2018 johannesUpdateDAQPSL acromag server moved from megatron to c1auxex2

I moved the epics IOC server process for the single Acromag ADC that monitors the PSL signals from megatron to c1auxex2.

First, I disabled the legacy support on all channels as explained in elog 13565. Then I copied the files npro_config.cmd and NPRO.db from /opt/rtcds/caltech/c1/scripts/Acromag to /cvs/cds/caltech/target/c1psl2/ following the pattern of the old Motorola machines and the new c1auxex2. I had to make some edits for correct paths and expanded the epics records to the standard we're using for ETMX.

I then added a service to systemd on c1auxex2 that runs the epics IOC for the Acromag PSL channels: /etc/systemd/system/modbusPSL.service. No more tmux on megatron.

Running two IOCs on a signle machine at the same time did not produce any errors and seems fine so far.

  13589   Wed Jan 31 15:27:55 2018 gautamUpdateSUSSUS MEDM master screens updated

I've often gotten confused by the labeling on the SUS MEDM screens about the coil "Vmon" fields - they're labelled as "30 Hz HPF", and indeed this is one of the many readbacks available on the coil driver board. But the actual EPICS channel that is being displayed in this field is from the "EPICS VMON" monitor point on the coil driver board. It has a gain of 1/2, so the actual voltage going to the coil is twice the channel value. Today, I fixed the SUS master screen to avoid this confusion - new labeling is shown in Attachment #1.

Attachment 1: NewSUSmaster.png
NewSUSmaster.png
  13588   Tue Jan 30 10:22:20 2018 SteveUpdateVACRGA scan at day 175

pd80b rga scan at 175 day.  IFO pressure 7.3e-6 Torr-it

Condition: vacuum normal, annuloses not pumped. Rga turned on yesterday.

The rga was not on since last poweroutage Jan 2, 2018 It is warming up and outgassing Atm2

Attachment 1: RGApd80-b-468-175d.png
RGApd80-b-468-175d.png
Attachment 2: RgaOn20hrs.png
RgaOn20hrs.png
  13587   Fri Jan 26 20:03:09 2018 gautamUpdateALSFiber ALS assay

I think part of the problem was that the rejected beam from the PBS was not really very Gaussian - looking at the spot on the beam profiler, I saw at least 3 local maxima in the intensity profile. So I'm now switching strategies to use a leakage beam from one of the PMC input steering optics- this isn't ideal as it already has the PMC modulation sideband on it, and this field won't be attenuated by the PMC transmission - but at least we can use a pre-doubler pickoff. This beam looks beautifully Gaussian with the beam profiler. Pics to follow shortly...

Quote:

I tried to couple the PSL pickoff into the fiber today for several hours, but got nowhere really, achieved a maximum coupling efficiency of ~10%. TBC tomorrow... Work done yesterday and today

 

  13586   Thu Jan 25 23:59:14 2018 gautamUpdateALSFiber ALS assay

I tried to couple the PSL pickoff into the fiber today for several hours, but got nowhere really, achieved a maximum coupling efficiency of ~10%. TBC tomorrow... Work done yesterday and today:

  • I changed the collimator from the fixed focal length but adjustable lens position CFC-2X-C to the truly fixed F220-APC-1064 recommended by johannes.
  • Used a pair of irises to level the beam out at 4" with two steering mirrors.
  • Used a connector on the PSL table to couple the EX laser light to the PSL fiber - then measured the mode using the beam-scanner (beam is ~300uW) 
  • Measured the mode of the PSL pickoff beam, also using the beam scanner. 
  • Per specs on the Thorlabs website, the F220-APC-1064 has a divergence angle of 0.032 degrees. So expected waist is ~1200um, and the Rayleigh range is ~4.3m, so this is not a very easy beam to measure and fit. I may be thinking about this wrong?
  • Measured beam 1/e^2 dia over ~0.65m, and found it to be fairly constant around 1800um (so waist of 900um) - beam is also pretty symmetric in x and y directions, but I didn't attempt an M^2 measurement.
  • The pickoff from the PSL also did not yield a very clean beam profile measurement, even though I measured over ~1m z-propagation distance. Nevertheless, this looked more like a Gaussian beam, and I confirmed the fitted waist size/location approximately by placing the beam profiler at the predicted waist location and checking the spot size.
  • Used jammt to calculate a candidate mode-matching solution - the best option seemed to be to use a combination of a f=150mm and f=-75mm lens in front of the collimator. 
  • Despite my best efforts, I couldn't get more than ~500uW of light coupled into the fiber - out of the 8mW available, this is a paltry 12.5% sad
  • Because the mode coming out of the fiber is relatively large, and because I have tons of space available on the PSL table, this shouldn't be a hard mode-matching problem, should be doable without any fast lenses - perhaps I'm doing something stupid and not realizing it. I'm giving up for tonight and will try a fresh assault tomorrow. 
  13585   Thu Jan 25 16:39:09 2018 KiraUpdatePEMSeismometer can insulation test

The final temperature reached in about 4.5 hours is 30.5C, while the starting temperature is about 24C. I can't seem to screenshot the data for some reason.

Also, I will calibrate the lab temperature sensor to Celcius in the near future so that we would have a working sensor inside the lab.

Quote:

After almost 3 hours the temperature rose by about 3.5C. Seems a bit slow, but we can drive it more if necssary. The heating curve itself is exponentiial, which is a good sign.

Quote:

We started the actual heating test today and it seems to be working so far. Hoping to heat it to about 40C. We also set up another temperature sensor to measure the lab temperature and connected it to J7, bottom.

 

 

Attachment 1: HeaterTest.png
HeaterTest.png
  13584   Thu Jan 25 14:58:20 2018 KiraUpdatePEMSeismometer can insulation test

After almost 3 hours the temperature rose by about 3.5C. Seems a bit slow, but we can drive it more if necssary. The heating curve itself is exponentiial, which is a good sign.

Quote:

We started the actual heating test today and it seems to be working so far. Hoping to heat it to about 40C. We also set up another temperature sensor to measure the lab temperature and connected it to J7, bottom.

 

  13583   Thu Jan 25 13:18:41 2018 gautamUpdateALSFiber ALS assay

I was looking at this a little more closely. As I understand it, the purpose of the audio differential IF amplifier is:

  1. To provide desired amplification at DC-audio frequencies
  2. To low pass the 2f component of the mixer output

Attachment #1 shows, the changes to the TF of this stage as a result of changing R19->50ohm, R17->500ohm. For the ALS application, we expect the beat signal to be in the range 20-100MHz, so the 2f frequency component of the mixer output will be between 40-200MHz, where the proposed change preserves >50dB attenuation. The Q of the ~500kHz resonance because of the series LCR at the input is increased as a result of reducing R17, so we have slightly more gain there. 

At the meeting yesterday, Koji suggested incorporating some whitening in the preamp itself, but I don't see a non-hacky way to use the existing PCB footprint and just replace components to get whitening at audio frequencies. I'm going to try and measure the spectrum of the I and Q demodulated outputs with the actual beat signal to see if the lack of whitening is going to limit the ALS noise in some frequency band of interest.

Does this look okay?

Quote:

The demod circuit board is configured to have gain of x100 post demod (conversion loss of the mixer is ~-8dB). This works well for the PDH cavity locking type of demod scheme, where the loop squishes the error signal in lock, so most of the time, the RF signal is tiny, and so a gain of x100 is good. For ALS, the application needs are rather different. So we lowered the gain of the "Audio IF amplifier" stage of the circuit from x100 to x10, by effecting the resistor swaps 10ohms->50ohms, 1kohm->500ohms (more details about this later).

Attachment 1: preampProposed.pdf
preampProposed.pdf
  13582   Thu Jan 25 12:41:18 2018 KiraUpdatePEMSeismometer can insulation test

We started the actual heating test today and it seems to be working so far. Hoping to heat it to about 40C. We also set up another temperature sensor to measure the lab temperature and connected it to J7, bottom.

Quote:

Gautam and I set up the insulated seismometer can in the lab today. I had previously wired up the two heaters I placed onto the sides of the can in parallel to get a total resistance of 12.5 ohms and then I wrapped the whole can in 3 layers of insulation (k-factor 0.25). We placed it on a large sheet of insulation as to not crush the wires leading out the bottom of the can. I stuck on one of my AD590 sensors to the inside of the can onto the copper lining using duct tape, though this is only a temporary solution. In the future, it would be nice to have some sort of thermal clamp to secure the sensor to the can. To provide power to the heater circuit board and the temperature sensor board, we got a powerstrip and plugged in two power supplies and a function generator into it. The heater circuit (attachment 3) is powered by one of the power supplies and the function generator, while the temperature sensor (attachment 5) is stuck to the side of the can and is powered by the second power supply. The heater circuit's MOSFET (IRF640, attachment 4) is placed on a metal block and sandwiched between two more to make sure it doesn't move around. The temperature sensor is connected by a long BNC cable to the channels in attachment 6.


gautam: we plugged the BNC output of Kira's temperature sensor circuit to J7 on the AA input chassis in 1X2 - this corresponds to ADC1 input 12 in c1ioo. I then made a "PEM" namespace block inside the c1als model, and placed a single CDS filter module inside it (this can be used for calibration purposes). The filter module is named "C1:PEM-SEIS_EX_TEMP", and has the usual CDSfilt channels available. I DQ'ed the output of the filter module (@256 Hz, probably too high, but I'm holding off on a recompile for now). Recompilation and model restart of c1als went smoothly. 

2 bench power supplies are being used for this test, we can think of a more permanent solution later.

**25 Jan noon: Added another filter module, "C1:PEM-SEIS_EX_TEMP", to which Kira is hooking up a second temperature sensor, which will serve as a monitor of the "Ambient" lab temperature. Added DQ channel for the output of this filter module, fixed sampling to 32Hz. Compile and restart went smooth.

 

  13581   Thu Jan 25 08:27:25 2018 SteveUpdatePEMearthquakes

M4 local earthquake at 10:10 UTC    There is no sign of damage.

....here is an other one.........M5.8 Ferndale, CA at 16:40 UTC

 

Attachment 1: M4_local_eq.png
M4_local_eq.png
Attachment 2: M4.png
M4.png
Attachment 3: M5.8Ferndale_CA.png
M5.8Ferndale_CA.png
  13580   Wed Jan 24 23:13:30 2018 johannesUpdateDAQc1auxex2 startup behavior
Quote:

The result is a smooth transition from idling to the controlled state with no sudden or large offset changes. yes

[Gautam, Johannes]

While checking how smooth the transition is we still noticed significant motion of ETMX by looking at the locked green laser and OpLevs. We found that this motion was not caused by interruption of the slow offset adjust, but rather the Watchdog being re-initialized to its OFF state, which cuts the fast channels OFF. On other optics this is observed too, but not as severe. The cause is a rather large offset on the LR coil coming from the fast DAQ, which was reported as 50mV by the slow readback channel (while other readback values are <10mV). It is present even when turning the output of the CDS model OFF, but vanishes when the watchdog is triggered. This helped us trace it to an offset of the DAC output itself: it is present at the output of the AI board but vanishes when the DAC is disconnected. The actual offset is ~40mV, as opposed to other channels on the same board, which ahve offsets in the range 3-7mV.

While we can compensate for this offset in software - it made us  wonder if the DAC channel is somehow busted and if that's what causing the 'wandering' of ETMX that we have been observing recently. There are two free DAC channels on the AI chassis that has the side coil and the green temperature control signals. We could re-route the LR signal through a different DAC channel to fix this.

gautam: 40mV offset at the AI board output gets multiplied by 3 in the dewhitening board, so there is a 120mV DC offset going to the coil (measured at dewhite board output with DMM). The offset itself isn't hurting us, but the fact that it is several times larger than other channels led us to wonder if it could be drifting around as well. From my SOS pitch balancing forays, in my head I have the number 30mrad as being the full range of the OSEM actuation - so if the offset swings by 120mV, that's ~150urad of motion, which is quite large, and is of the order of magnitude I'm used to seeing ETMX move around by.

  13578   Wed Jan 24 19:17:06 2018 johannesUpdateDAQc1auxex2 startup behavior

I compiled the burt binaries on c1auxex2 which took a little fiddling with dependencies and paths but nothing too major. The complete local epics folder (/opt/epics/) which contains the base epics binaries, modbus and burt for 32-bit linux has been copied to the shared drive at /opt/rtapps/epics-3.15.5. They belong to the most recent stable release. This was so we can now automatically call burt after the IOC initialization on c1auxex2 to restore the backed-up channel values.

I also copied the database definition and modbus instruction files to /cvs/cds/caltech/target/c1auxex2, from where they are now being read upon IOC initialization. This is an excerpt of the service file:

#ExecStart=/usr/bin/procServ -f -L /home/controls/modbusIOC/modbusIOC.log -p /run/modbusioc.pid 8008 /opt/epics/modules/modbus/bin/linux-x86/modbusApp /cvs/cds/caltech/target/c1auxex2/ETMXaux2.cmd   <-- Contains logging to file, see note 1)
ExecStart=/usr/bin/procServ -f -p /run/modbusioc.pid 8008 /opt/epics/modules/modbus/bin/linux-x86/modbusApp /cvs/cds/caltech/target/c1auxex2/ETMXaux2.cmd <-- Initializes the EPICS IOC with Modbus support
ExecStop=/bin/kill -9 ` cat /run/modbusioc.pid` <-- Kills the detached process by its process ID
ExecStartPost=/bin/bash -c "/opt/epics/extensions/bin/linux-x86/burtwb -f /opt/rtcds/caltech/c1/burt/autoburt/latest/c1auxex.snap" <-- Restores general channel values
ExecStartPost=/bin/bash -c "/opt/epics/extensions/bin/linux-x86/burtwb -f /opt/rtcds/caltech/c1/medm/MISC/ifoalign/burt/ETMX.snap" <-- Restores PIT and YAW values from align MEDM screen
ExecStartPost=/bin/bash -c ". /home/controls/modbusIOC/ETMXaux2.sh" <-- Enables writing to PIT and YAW DAC channels, see note 2)

Note 1) I removed the logging to file for now because I noticed that if there are Acromag communication issues the logfile tends to grow in size VERY fast. In the cryo lab is had gotten to over 70GB just over the winter break. I don't think it's absolutely necessary to have it, and if diagnostics are needed we can easily uncomment it temporarily.

Note 2) I modified the static EPICS records of the four OSEM bias adjust channels so they won't start updating as soon as the IOC starts up (and before the channel defaults are restored by burt). This was done by setting the OMSL (output mode select) field from "closed_loop" to "supervisory". Sample record:

record(ao,"C1:SUS-ETMX_ULBiasAdj")
{
        field(DESC,"Bias Adjust for ETMX UL Coil Output")
        field(DTYP,"asynInt32")
        field(OUT, "@asynMask(C1AUXEX_XT1541A_DAC, 0, -16)MODBUS_DATA")
        field(SCAN,".1 second")
        field(OMSL,"supervisory")  <-- Used to be "closed_loop"
        field(DOL, "C1:SUS-ETMX_ULBiasSet  PP")
        field(PREC,"3")
        field(EGUF,"10.923")
        field(EGUL,"-10.923")
        field(EGU, "Volts")
        field(LINR,"LINEAR")
        field(DRVH,"10")
        field(DRVL,"-10")
        field(HOPR,"10")
        field(LOPR,"-10")
}

Now, on reboort/IOC re-initialization the physical DAC channels are performing a one-time readback of the last stored value in the Acromag's register, then idle until the last StartPost statement executes the script ETMXaux.sh, which changes their OMSL field back to "closed_loop". This causes them to start updating their output from the calc records defined in their DOL field (which have by then recovered their default values curtesy of burt). The result is a smooth transition from idling to the controlled state with no sudden or large offset changes. yes

  13576   Wed Jan 24 18:12:31 2018 johannesUpdateDAQETMX auxiliary DAQ work

I replaced the two remaining D-Sub M/M cables that still had gender-changers with M/F cables today, completing the mechanical and wiring work on the ETMX rack. All backplane adapter boards were secured to a cross-strut of the crate using zip ties. This was necessary because the adapter boards don't fit the crate with their panels attached ( the ETMX eurocrate is the only one with slightly different dimensions from all the others), and the we can't mount them to the strut using the panels. This won't be an issue on any of the other crates.

   

   

In other news:

I disabled the legacy support in the three Acromag ADC units and set the input averaging to 10 samples via the USB configuration utility. The documentation is unfortunately a little sparse about what this actually means. The manual states that "fresh input data is available to the network every 10ms", so the sampling rate is for sure faster than 100Hz. Since the IOC updates its channels every .1 seconds I assume that an averaging value of 10 to reduce the input noise is safe. The maximum value the configuration tool permits is 200. I tried this using the CryoLab DAQ and set all input channels to 200 and used StripTool to look at the time series of a slow oscillation (.1Hz) with a large amplitude (16Vpp) and looked for missed data points, indicating too long wait times for channels updates. There was no such qualitative difference between 1 sample, 10 samples, and 200 samples, so even pushing the averaging value to max seemed okay. I went with the conservative value of 10 for the ETMX DAQ, but we can likely increase this if noise on the slow inputs becomes an issue.

The input scaling of the ADC channels has been corrected. I changed the conversion method in the EPICS records from manual using the ASLO and AOFF fields to using engineering units via EGUF and EGUL. This required a little attention. The Acromags scale the dynamic input range of +/- 10V to +/- 30,000 raw value, but the EPICS IOC interprets the data type as ranging from -32767 to +32768, so the EGUF and EGUL fields must be set to -10.923 and +10.923 to achieve proper scaling. I also changed the SCAN field on all ADC channels to 0.1 seconds. This has been done for all ADC and DAC channel records.

  13575   Wed Jan 24 13:55:04 2018 KiraSummaryPEMSeismometer can insulation test

Gautam and I set up the insulated seismometer can in the lab today. I had previously wired up the two heaters I placed onto the sides of the can in parallel to get a total resistance of 12.5 ohms and then I wrapped the whole can in 3 layers of insulation (k-factor 0.25). We placed it on a large sheet of insulation as to not crush the wires leading out the bottom of the can. I stuck on one of my AD590 sensors to the inside of the can onto the copper lining using duct tape, though this is only a temporary solution. In the future, it would be nice to have some sort of thermal clamp to secure the sensor to the can. To provide power to the heater circuit board and the temperature sensor board, we got a powerstrip and plugged in two power supplies and a function generator into it. The heater circuit (attachment 3) is powered by one of the power supplies and the function generator, while the temperature sensor (attachment 5) is stuck to the side of the can and is powered by the second power supply. The heater circuit's MOSFET (IRF640, attachment 4) is placed on a metal block and sandwiched between two more to make sure it doesn't move around. The temperature sensor is connected by a long BNC cable to the channels in attachment 6.


gautam: we plugged the BNC output of Kira's temperature sensor circuit to J7 on the AA input chassis in 1X2 - this corresponds to ADC1 input 12 in c1ioo. I then made a "PEM" namespace block inside the c1als model, and placed a single CDS filter module inside it (this can be used for calibration purposes). The filter module is named "C1:PEM-SEIS_EX_TEMP", and has the usual CDSfilt channels available. I DQ'ed the output of the filter module (@256 Hz, probably too high, but I'm holding off on a recompile for now). Recompilation and model restart of c1als went smoothly. 

2 bench power supplies are being used for this test, we can think of a more permanent solution later.

**25 Jan noon: Added another filter module, "C1:PEM-SEIS_EX_TEMP", to which Kira is hooking up a second temperature sensor, which will serve as a monitor of the "Ambient" lab temperature. Added DQ channel for the output of this filter module, fixed sampling to 32Hz. Compile and restart went smooth.

Attachment 1: IMG_20180124_124209.jpg
IMG_20180124_124209.jpg
Attachment 2: IMG_20180124_124202.jpg
IMG_20180124_124202.jpg
Attachment 3: IMG_20180124_124224.jpg
IMG_20180124_124224.jpg
Attachment 4: IMG_20180124_124229.jpg
IMG_20180124_124229.jpg
Attachment 5: IMG_20180124_124236.jpg
IMG_20180124_124236.jpg
Attachment 6: IMG_20180124_124156.jpg
IMG_20180124_124156.jpg
  13574   Wed Jan 24 10:45:14 2018 gautamUpdateALSFiber ALS assay

I was looking into the physics of polarization maintaining fibers, and then I was trying to remember whether the fibers we use are actually polarization maintaining. Looking up the photos I put in the elog of the fibers when I cleaned them some months ago, at least the short length of fiber attached to the PD doesn't show any stress elements that I did see in the Thorlabs fibers. I'm pretty sure the fiber beam splitters also don't have any stress elements (see Attached photo). So at least ~1m of fiber length before the PD sensing element is probably not PM - just something to keep in mind when thinking about mode overlap and how much beat we actually get.  

 

  13573   Wed Jan 24 00:58:59 2018 gautamUpdateALSX Green PDH modulation depth

On Friday, while Udit and I were doing some characterization of the EX+PSL IR beat at the LSC rack, I noticed that there were sidebands around the main beat peak at 20dBm lower level. These were offset from the main peak by ~200kHz - I didn't do a careful characterization but because of the symmetric nature of these sidebands and the fact that they appeared with the same offset from the main peak for various values of the central beat frequency, I hypothesize that these are from the modulation sidebands we use for PDH locking the EX laser to the arm cavity. So we can estimate the modulation depth from the relative powers of the main beat peak and the ~200kHz offset sidebands.

Since the IR light is used for the beat and we directly couple it to the fiber to make the beat, there is no green or IR cavity pole involved here. 20dBm in power means \frac{\beta^2}{4} \approx 10^{\frac{-20}{10}} \approx 0.01. And so the modulation depth, \beta \approx 0.2 \mathrm{rad}. I will do a more careful meaurement of this, but this method of measuring the modulation depth can give us a precise estimate - for what it's worth, this number is in the same ballpark as the measurement I quote in elog12105.

What is the implication of having these sidebands on our ALS noise? I need to think about this, effectively the phase noise of the SR function generators we use to do the phase modulation of the EX laser is getting imprinted on the ALS noise? Is this hurting us in any frequency range that matters?

 

  13572   Wed Jan 24 00:48:47 2018 gautamUpdateElectronicsWhitening filter D990694

I plan to do some characterization of this problem. The plan is to use THD as a metric for whether we are having hidden saturations. Pg 9 of the LT1125 datasheet tells us what fraction of THD to expect. I will use one of the several unused DAC channels available at the LSC rack to drive a 100Hz sine wave into one of the inputs of the whitening chassis, and measure the THD up to a reasonable harmonic number (will probably be set by the ADC noise) for (i) various whitening gain settings and (ii) various input signal amplitudes.

The motivation is to attempt to quantify the problem better:

  1. How bad is it to have one or more of the OpAmps in the quad IC either saturated to its voltage rails, or max output current?
  2. Can we reproduce Hartmut's observations?
  3. Are the OpAmps already irreversibly damaged because of extended abuse?

Then we can decide what, if anything, to do about this issue.

  13571   Wed Jan 24 00:33:31 2018 gautamUpdateALSFiber ALS assay

I did some work on the PSL table today. Main motivations were to get a pickoff for the BeatMouth PSL beam before any RF modulations are imposed on it, and to improve the mode-matching into the fiber. Currently, we use the IR light reflected by the post doubling oven harmonic separator. This has the PMC modulation sideband on it, and also some green leakage. 

So I picked off ~8.5mW of PSL light from the first PBS (pre Faraday rotator), out of the ~40 mW available here, using a BS-80-1064-S. I dumped the 80% reflected light into the large beam dump that was previously being used to dump this PBS reflection. Initially, I used a R=10% BS for S-pol that I found on the SP table, but Koji tipped me off on the fact that these produce multiple reflected beams, so I changed strategy to use the R=80% BS instead.

The transmitted 20% is routed to the West edge of the PSL table via 2 1" Y1-1037-45S optics, towards the rough vicinity of the fiber coupler. For now it is just dumped, tomorrow I will work on the mode matching. We may want to cut the power further - ideally, we want ~2.5mW of power in the fiber - this is then divided by 4 inside the beat mouth before reaching the beat PD, and with other losses, I expect ~500mW of PSL power and comparable AUX light, we will have a strong >0dBm beat.

Attachment #1 is a picture of my modifications. For this work, I

  • Closed PSL shutter, turned HEPA up 
  • Moved HP GHz spec analyzer to the side for ease of access to the table.
  • Moved several optics that look to me as to have once been part of the RefCav setup - I don't think this would have been a useful alignment reference in any case as we moved the RefCav in a non-deterministic way for the PSL secondary shelf install.
  • Used one 1" 45 deg S-pol optic from the optics cabinet - remaining optics were scavenged from PSL table and SP table.
  • Removed an SMA cable connected to an EOM, whose other end wasn't connected to anything.
  • Turned HEPA back down, IMC locks fine now.

 

Attachment 1: IMG_6866.JPG
IMG_6866.JPG
  13570   Tue Jan 23 16:02:05 2018 SteveConfigurationSEIload cells

1500 and 2000 lbs load cells arrived from MIT to measure the vertical loads on each leg.

Quote:

We've been thinking about putting in a blade spring / wire based aluminum breadboard on top of the ETM & ITM stacks to get an extra factor of 10 in seismic attenuation.

Today Koji and I wondered about whether we could instead put something on the outside of the chambers. We have frozen the STACIS system because it produces a lot of excess noise below 1 Hz while isolating in the 5-50 Hz band.

But there is a small gap between the STACIS and the blue crossbeams that attache to the beams that go into the vacuum to support the stack. One possibility is to put in a small compliant piece in there to gives us some isolation in the 10-30 Hz band where we are using up a lot of the control range. The SLM series mounts from Barry Controls seems to do the trick. Depending on the load, we can get a 3-4 Hz resonant frequency.

Steve, can you please figure out how to measure what the vertical load is on each of the STACIS?

 

Attachment 1: stacis3LoadCells.png
stacis3LoadCells.png
Attachment 2: DSC00025.JPG
DSC00025.JPG
Attachment 3: DSC00026.JPG
DSC00026.JPG
  13569   Tue Jan 23 01:56:18 2018 gautamUpdateElectronicsTeledyne AP1053

I have acquired 5 pieces of the Teledyne AP1053 from Koji - these are now at the 40m. I will determine an appropriate location for storage of these and update. We are also looking to acquire 5 more of these. The combination of high power output (26dBm), low gain (10dB), and low noise figure (1.5dB) are quite uncommon in an amplifier and so these should be used only when such properties are required simultaneously.

*Steve informs me that these amps have been stored in the RF cabinet E6 along the east arm.

Steve's note: Teledyne rf amp product selection guide

                   Teledyne rf low noise amp guide

 

Attachment 1: DSC00022.JPG
DSC00022.JPG
  13568   Tue Jan 23 01:33:23 2018 gautamUpdateElectronicsWhitening filter D990694

After discussing with Koji, we looked at the aLIGO incarnation of this board. Interestingly, it too has a similar topology of 4 switchable gain stages with gains of 24, 12, 6 and 3dB. The main differences are that they use single Op27 ICs instead of the quad LT1125s, and also, they use a different combination of feedback resistors to realize the various gains. 

We considered upping the feedback resistance (R15, R143) on the 24dB gain stage of our boards from (1k, 66.5ohms) to (3k, 200ohms) as on the aLIGO boards - but this doesn't really help? Because KCL demands that the same current flow in R15 and R143, and so the output Vsat of the op amp and its max current driving capabilities in combination determine if the inverting input can follow the non inverting input?

As Hartmut points out in his note, he was able to access the full range of ADC voltages when the gain was set to 3dB, despite the fact that the LT1125 was still getting internally saturated. Operating with minimum 24dB whitening gain doesn't really solve the problem either because the problem just gets shifted to the next gain stage in the chain, and we still have saturation. I also don't have a feeling for how much differential voltage these LT1125s can sustain before they are damaged - I guess the planned THD check will reveal if they are okay or not.

It seems to me like the only way to truly fix this problem of one stage saturating and screwing up the others is to use single Op27s (or equivalent) in place of the quad LT1125s. The aLIGO design also has a series resistance to the non-inverting input - this can help prevent current overdraw from the previous stage (due to a lowered input impedance of the OpAmp - but I wonder how low this can go?).

Quote:

this is the note from Hartmut Grote on this topic from 2004

 

  13567   Mon Jan 22 20:54:58 2018 KojiSummaryGeneralAUX-PSL beat setup

The beat setup has been made on the PSL table. The BS and the PD were setup. The beat was found at 29.42degC and 50.58degC for the PSL and AUX crystal temperatures, respectively.
We are ready for the EOM test. I have instruments stacked around the PSL table. Please leave them as they are for a while. If you need to move them, please contact with me. Thanks.


A picked-off PSL after the main modulator was used as the PSL beam. This was already introduced close to the setup thanks to the previous 3f cancellation test ELOG 11029. The AUX beam was obtained from the transmission of 90% mirror. Both paths have S polarization. The beams are combined with a S-pol 50% BS. The combined beam is detected by a new focus 1GHz PD.

The PSL crystal temp (actual) was 50.58degC. The AUX crystal temp was swept upward and the string beat was found at 50.58degC. After a bit of alignment, the beat strength was -18dBm (at 700V/A RF transimpedance of NF1611) .

Attachment 1: IMG_3051.JPG
IMG_3051.JPG
  13566   Mon Jan 22 12:48:48 2018 KojiSummaryGeneralBeat setup for aLIGO EOM test

I'm planning to construct a beat setup between the PSL and AUX beams. I am going to make it in the area shown in a blue square in the attached photo. This does not disturb Johannes' and PSL setups. The beams are obtained from the PBS reflection of the PSL and the dumped beam of the aux path (0th or 1st order beam of the AOM).

Attachment 1: IMG_3048.JPG
IMG_3048.JPG
  13565   Sun Jan 21 13:11:25 2018 johannesUpdateDAQAcromag checks

After some research: -the- reason for the reduced +/- 20,000 swing in raw values is a default setting for having support for legacy devices enabled when using the acromag proprietary i2o peer-to-peer protocol. So this is doubly unnecessary because a) we don't have any legacy devices at all and b) we're using pure modbus/TCP and no i2o. To change the setting I have to connect via the USB configuration utility. In addition, I want to understand the averaging feature of the acromag units better, which is also configured via USB, and lets one set a fixed amount of samples to be averaged before updating the read-register value. The documentation says that the 8 channels are multiplexed into a single ADC, and that new input data is available after 10 ms for each channel, suggesting a sampling rate of 100 Hz per channel and that the multiplexing happens faster, but is not super-clear about this, so I want to test it in the cryo lab first before unleashing it onto c1auxex2.

Furthermore, the standard timing options for updating epics records are 10s, 5s, 2s, 1s, 0.5s, 0,2s, and 0.1s. On the previous c1auxex, the monitoring channels were set to 0.1s, but that clashes with the 16 Hz global EPICS rate, resulting in partial double-sampling. One can manually provide the option 0.0625s for 16Hz update rate. I will test this and how it deals with the averaging in the cryolab too.

  13564   Sat Jan 20 15:57:11 2018 ranaUpdateElectronicsWhitening filter D990694

this is the note from Hartmut Grote on this topic from 2004

  13563   Sat Jan 20 01:20:37 2018 gautamUpdateElectronicsWhitening filter D990694

We use D990694 in various places. Today, Rana alerted me to an important consideration to be kept in mind when we use this board, which I found quite interesting. I still don't understand the problem at the BJT level, but I think one can appreciate the problem without going to the transistor design of the LT1125. I'm attaching an annotated schematic of the whitening section in question. If the following assumptions are valid, then I think my picture is valid.

  1. The switch used to bypass the various whitening gain stages, namely the ADG333ABR, has infinite impedance in the "OFF" state, such that when the 24dB gain stage is bypassed, U28A (or in general one of the 4 quad op-amps) is forced to drive it's output voltage across 1.0665 kohms of resistance.
  2. The individual LT1125 Op Amps can drive a maximum of 30mA of current.

Then, as one can see in the attached schematic, when we set the gain of any input to <24dB, we must ensure that the input voltage is less than approximately 2V. Otherwise, by asking too much of the first stage op-amp in the quad IC LT1125, we may be messign around with all the 4 op amps in the quad! Even the 0dB setting is not immune to this problem, as it uses one of the 4 op amps.

I don't think the usual rules of calculating the gain of a non-inverting amplifier (G = 1 + Rf/Ri) remain valid even when the op-amp is forced to drive more output current than it can, and I don't have a way to quantify the possible interference between the 4 op amps in the quad - but does this seem like a valid conclusion? If so, we must check signal levels of various LSC signals. AS55 signals currently have the 0dB gain setting - I had turned this down from 6dB some months ago, because it seemed like the ADC was saturating at the 6dB gain setting, which suggests that the input voltage is ceratinly > 2V, and AS55_Q is what is used for MICH control in the DRMI. All of my noise budgeting work over the last few months used this setting, I wonder if they are all invalid surprise


Now that I think about this a bit more - this problem shouldn't be significant for the usual LSC degrees of freedom when in lock, as the huge DC gain of the loop should squish large DC values of the error signals, and so there shouldn't be any danger of overloading the LT1125. But I don't know if we are being hurt by this effect when flashing through resonances, when the PDH horn-to-horn voltage can be quite high (which is in principle a good thing?). I don't know if there is any "hysterisis" effect where the overloaded quad IC has some relaxation time before it returns to normal operation, and if we are being limited in our ability to catch lock because if this effect. 

The concerns remain valid for th ALS demodulated error signals though, for which the signals will remain large throughout.

Attachment 1: whiteningBoardLimitations.pdf
whiteningBoardLimitations.pdf
ELOG V3.1.3-